# A reanalysis of event shape distributions in electron-positron annihilation

### v1

@@ Abstract
I Introduction
@@
Event shape variables have been extensively studied at the Large Electron Positron (LEP) collider performing high precision measurements Abdallah:2003xz ; Heister:2003aj ; Achard:2004sv ; Abbiendi:2004qz ; SLD:1994idb . In order to describe experimental data with high accuracy, it is crucial to achieve the highest possible theoretical precision. Fixed order calculations for event shape observables have been performed up to the NNLO level of accuracy and are shown in Refs. GehrmannDeRidder:2007hr ; Gehrmann-DeRidder:2008qsl ; Weinzierl:2008iv ; Weinzierl:2009ms ; DelDuca:2016ily ; DelDuca:2016csb . Results including resummation of infra-red large logarithms have also been performed in Refs. Banfi:2004nk ; Banfi:2004yd ; Banfi:2014sua ; Banfi:2016zlc ; Chiu:2011qc ; Chiu:2012ir ; Chien:2010kc ; Becher:2012qc ; Abbate:2010xh ; Hoang:2014wka .
@@
The remaining sections of this paper are organized as follows. We present in Sec. II the calculation technology for applying the PMC method to event shape observables in electron-positron annihilation; in Sec. III, we give theoretical predictions and discussions; Sec. IV is reserved for a summary.
II PMC scale-setting for event shape observables
@@
$\frac{1}{\sigma_{tot}}\frac{d\sigma}{dy}$ $=$ $\bar{A}(y,\mu_{r})^{V}\,a^{V}_{s}(\mu_{r})+\bar{B}(y,\mu_{r})^{V}\,a^{V,2}_{s}(\mu_{r})$ (7)
$+{\cal O}(a^{V,3}_{s}),$
@@
$\frac{1}{\sigma_{tot}}\frac{d\sigma}{dy}$ $=$ $\bar{A}(y,\mu_{r})^{V}\,a^{V}_{s}(Q_{*})+\bar{B}(y,\mu_{r})^{V}_{\rm con}\,a^{V,2}_{s}(Q_{*})$ (9)
$+{\cal O}(a^{V,3}_{s}),$
@@
III Theoretical predictions and discussions
@@
In this paper, we show results for the differential distributions of the following event shape variables: thrust ($T$), heavy jet mass ($\rho=M^{2}_{H}/s$), wide jet broadening ($B_{W}$), total jet broadening ($B_{T}$), C-parameter ($C$) and jet-transition ($Y_{3}$). These event shape distributions have been measured at LEP by several experiments and at different center-of-mass energies ranging from $35$ to $206$ GeV Heister:2003aj .
III.1 Event shape distributions at the $Z^{0}$ peak using the conventional scale-setting method
@@
III.2 Event shape distributions at the $Z^{0}$ peak using the PMC scale-setting method
Figure 2: PMC scales at $\sqrt{s}=91.2$ GeV in the $V$ scheme (solid line) and in $\overline{\rm MS}$ scheme (dashed line) for the event shape variables: thrust ($1-T$), heavy jet mass ($\rho$), wide jet broadening ($B_{W}$), total jet broadening ($B_{T}$) and C-parameter ($C$) respectively.
@@ III.2 Event shape distributions at the $Z^{0}$ peak using the PMC scale-setting method
Besides, we also notice clear deviations from data of the five event shapes observables in both the two-jet and the multijet regions. These are due to the presence of large IR-logarithms that are enhanced in these regions spoiling the perturbative fixed-order QCD predictions. It is possible to restore the correct perturbative regime by resumming the logarithmically enhanced terms to all orders in the coupling constant. The resummation of large logarithms has been extensively studied in the literature. In fact, due to kinematical constraints, the event shape distribution at LO is limited to a certain range. For instance, the thrust distribution is restricted to the range of $30\leq(1-T)\leq 1/$ 3 and the $C$-parameter distribution is restricted to the range of $0\leq C\leq 0.75$. Hence, in this region, PMC predictions are dominated by the 3-jet process contributions and thus show a better agreement with the accurate experimental data. Outside this range, pQCD calculations are significantly affected by missing higher order contributions and thus the 3-jet process contribution alone cannot thoroughly describe data. As reported in literature, in order to obtain a correct analysis of event shape variables performed using the conventional method, one has to include also the resummation of large logarithms outside the kinematical constrained range. Furthermore, in addition to the resummation of large logarithms, non-perturbation effects should be considered in the two-jet region.
@@
III.3 Event shape distributions above the $Z^{0}$ peak and below the $Z^{0}$ peak using the conventional scale-setting method and the PMC scale-setting method
@@
Figure 6: Results for thrust distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=133$, $161$, $172$, $183$, $189$, $200$, $206$ GeV respectively.. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$, while error bands for the PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from ALEPH Heister:2003aj .
@@
Figure 7: Results for C-parameter ($C$) distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=133$, $161$, $172$, $183$, $189$, $200$, $206$ GeV respectively. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$, while error bands for the PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from ALEPH Heister:2003aj .
@@
In addition to the event shape distributions at the center-of-mass energy $\sqrt{s}=91.2$ GeV, we also present results for event shape distributions at higher energies. Results for thrust ($T$) and C-parameter ($C$) distributions are shown in Figs.(6) and (7) respectively, including the comparison between the conventional and the PMC results. Results for each event shape variable, at different energies, i.e. $\sqrt{s}=133$, $161$, $172$, $183$, $189$, $200$, $206$ GeV, are shown together with their experimental data measured with high accuracy at the LEP facility Abdallah:2003xz . In Figs.(6) and (7), we notice that the PMC results for both the thrust ($T$) and the C-parameter ($C$) have been improved at $\sqrt{s}=133$, $161$, $172$, $183$, $189$, $200$, $206$ GeV. Given the large uncertainties of the experimental data at $\sqrt{s}=161$, $172$ GeV, both the conventional and the PMC results lead to analogous predictions if compared to the experimental data. While at the energy of $\sqrt{s}=133$, $183$, $189$, $200$, $206$ GeV, PMC results have a better fit with the experimental data if compared to the conventional results.
@@
Figure 8: Results for thrust distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=35$, $44$ GeV respectively. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$. Error bands for PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from JADE MovillaFernandez:1997fr ; Biebel:1999zt .
@@
Figure 9: Results for the C-parameter ($C$) distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=35$, $44$ GeV respectively. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$. Error bands for PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from JADE MovillaFernandez:1997fr ; Biebel:1999zt .

### v2

@@ Abstract
Keywords:
Jets, Perturbative QCD, Renormalization scale and scheme, Renormalization group equation
1 Introduction
@@
Event shape variables have been extensively studied at the Large Electron Positron (LEP) collider performing high precision measurements Abdallah:2003xz ; Heister:2003aj ; Achard:2004sv ; Abbiendi:2004qz ; SLD:1994idb . In order to describe experimental data with high accuracy, it is crucial to achieve the highest possible theoretical precision. Fixed order calculations for event shape observables have been performed up to the NNLO level of accuracy and are shown in Refs. GehrmannDeRidder:2007hr ; Gehrmann-DeRidder:2008qsl ; Weinzierl:2008iv ; Weinzierl:2009ms ; DelDuca:2016ily ; DelDuca:2016csb ; Gehrmann:2017xfb . Results including resummation of infra-red large logarithms have also been performed in Refs. Banfi:2004nk ; Banfi:2004yd ; Banfi:2014sua ; Banfi:2016zlc ; Chiu:2011qc ; Chiu:2012ir ; Chien:2010kc ; Becher:2012qc ; Abbate:2010xh ; Hoang:2014wka ; Abbate:2012jh ; Hoang:2015hka ; Benitez:2024nav ; Benitez:2025vsp .
@@
The remaining sections of this paper are organized as follows. We present in Sec. 2 the calculation technology for applying the PMC method to event shape observables in electron-positron annihilation; in Sec. 3, we give theoretical predictions and discussions; Sec. 4 is reserved for a summary.
2 PMC scale-setting for event shape distributions
@@
$\frac{1}{\sigma_{tot}}\frac{d\sigma}{dy}=\bar{A}(y,\mu_{r})^{V}\,a^{V}_{s}(\mu_{r})+\bar{B}(y,\mu_{r})^{V}\,a^{V,2}_{s}(\mu_{r})+{\cal O}(a^{V,3}_{s}),$ (7)
@@
$\frac{1}{\sigma_{tot}}\frac{d\sigma}{dy}=\bar{A}(y,\mu_{r})^{V}\,a^{V}_{s}(Q_{*})+\bar{B}(y,\mu_{r})^{V}_{\rm con}\,a^{V,2}_{s}(Q_{*})+{\cal O}(a^{V,3}_{s}),$ (9)
@@
3 Theoretical predictions and discussions
@@
In this paper, we show results for the differential distributions of the following event shape variables: thrust ($T$), heavy jet mass ($\rho=M^{2}_{H}/s$), wide jet broadening ($B_{W}$), total jet broadening ($B_{T}$), C-parameter ($C$) and jet-transition ($Y_{3}$). These event shape distributions have been measured at LEP by several experiments and at different center-of-mass energies ranging from 353535 to 206206206 GeV Heister:2003aj .
3.1 Event shape distributions at the $Z^{0}$ peak using conventional scale setting
@@
3.2 Event shape distributions at the $Z^{0}$ peak using PMC scale setting
[…]
@@ 3.2 Event shape distributions at the $Z^{0}$ peak using PMC scale setting
The soft-limit of the PMC in the two-jet region is not due to an ad-hoc introduction of a particular kinematic cutoff in the PMC calculation, but it is a direct consequence of the $N_{f}$-terms that stem from the loop-integration and thus it is directly connected to the loop kinematic boundary. In fact the PMC-scale dynamics reflects the dynamics of the virtuality of the gluons and quarks in the loops, this virtuality is constrained by kinematics and thus it leads to the zero-limit in the two-jet region. We remind that PMC methodology applies to the perturbative QCD calculation and thus the validity range of the PMC-scale is determined by the $Q_{*}\gg\Lambda_{\rm QCD}$ condition that is verified in the following bin-range of each variable at $\sqrt{s}=91.2$ GeV: $0.01\leq(1-T)\leq 0.33$, $0.02\leq\rho\leq 0.33$, $0.04\leq B_{W}\leq 0.29$, $0.04\leq B_{T}\leq 0.29$, $0.06\leq C\leq 0.75$ where the PMC-scale is: $Q_{*}>1.5$ GeV and thus in the perturbative region.
@@
Besides, we also notice clear deviations from data of the five event shapes observables in both the two-jet and the multijet regions. These are due to the presence of large IR-logarithms that are enhanced in these regions spoiling the perturbative fixed-order QCD predictions. It is possible to restore the correct perturbative regime by resumming the logarithmically enhanced terms to all orders in the coupling constant. The resummation of large logarithms has been extensively studied in the literature. In fact, due to kinematical constraints, the event shape distribution at LO is limited to a certain range. For instance, the thrust distribution is restricted to the range of $0\leq(1-T)\leq 1/3$ and the $C$-parameter distribution is restricted to the range of $0\leq C\leq 0.75$. Hence, in this region, PMC predictions are dominated by the 3-jet process contributions and thus show a better agreement with the accurate experimental data. Outside this range, pQCD calculations are significantly affected by missing higher order contributions and thus the 3-jet process contribution alone cannot thoroughly describe data. As reported in literature, in order to obtain a correct analysis of event shape variables performed using the conventional method, one has to include also the resummation of large logarithms outside the kinematical constrained range. Furthermore, in addition to the resummation of large logarithms, non-perturbation effects should be considered in the two-jet region.
We remark that a partial resummation of the IR large-logarithms already occur in the perturbative calculations up to NNLO. In fact, this explains the presence of the peaks in the distributions and in the two-jet region. This partial resummation is not related to the PMC, but is simply due to the improvement of the precision with the order of accuracy and the occurrence of the IR-large-logarithms at higher order resum with those of the lower orders. This partial resummation is incomplete, since the series is truncated at the NNLO. Thus, it cannot restore the correct behavior of the distributions in the region on left-side of the peak close to the left-boundary, where instead the distributions oscillate severely. Thus, in this region the entire resummation procedure is required in order to restore the correct predictive behavior of the perturbative calculation. Besides, in Ref. DiGiustino:2021nep , it was also shown that, even though the presence of the peak is due to resummation effects, its position is more related to the renormalization scale and in particular to the PMC scale. Thus, the use of the PMC can improve the perturbative theoretical predictions, extending the range of validity of the fixed-order calculation and the comparison with data. The PMC scale itself can indicate the dynamical range of validity of the perturbative calculation and the region where also other effects such as the resummation of the IR-large-logarithms must be taken into account, in order to have a complete and correct description of the physics all over the entire physical range [^2] A more detailed study of the interplay between PMC and Resummation for event shape variables is in progress..
@@
3.3 Event shape distributions above and below the $Z^{0}$ peak
@@
Figure 6: Results for thrust distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=133$, 161161161, 172172172, 183183183, 189189189, 200200200, 206206206 GeV respectively.. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$, while error bands for the PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from ALEPH Heister:2003aj .
@@
Figure 7: Results for C-parameter ($C$) distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=133$, 161161161, 172172172, 183183183, 189189189, 200200200, 206206206 GeV respectively. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$, while error bands for the PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from ALEPH Heister:2003aj .
@@
In addition to the event shape distributions at the center-of-mass energy $\sqrt{s}=91.2$ GeV, we also present results for event shape distributions at higher energies. Results for thrust ($T$) and C-parameter ($C$) distributions are shown in Figs.(6) and (7) respectively, including the comparison between the conventional and the PMC results. Results for each event shape variable, at different energies, i.e. $\sqrt{s}=133$, 161161161, 172172172, 183183183, 189189189, 200200200, 206206206 GeV, are shown together with their experimental data measured with high accuracy at the LEP facility Abdallah:2003xz . In Figs.(6) and (7), we notice that the PMC results for both the thrust ($T$) and the C-parameter ($C$) have been improved at $\sqrt{s}=133$, 161161161, 172172172, 183183183, 189189189, 200200200, 206206206 GeV. Given the large uncertainties of the experimental data at $\sqrt{s}=161$, 172172172 GeV, both the conventional and the PMC results lead to analogous predictions if compared to the experimental data. While at the energy of $\sqrt{s}=133$, 183183183, 189189189, 200200200, 206206206 GeV, PMC results have a better fit with the experimental data if compared to the conventional results.
@@
Figure 8: Results for thrust distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=35$, 444444 GeV respectively. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$. Error bands for PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from JADE MovillaFernandez:1997fr ; Biebel:1999zt .
@@
Figure 9: Results for the C-parameter ($C$) distribution using the conventional scale-setting (Conv.) at LO (dashed), NLO (dotdashed ) and NNLO (dotted) GehrmannDeRidder:2007hr and results for the PMC scale-setting (solid line) are shown for $\sqrt{s}=35$, 444444 GeV respectively. Error bands for the conventional predictions are obtained by varying $\mu_{r}\in[\sqrt{s}/2,2\sqrt{s}]$. Error bands for PMC predictions are given by the uncertainty $\Delta\alpha_{s}(M_{Z})=\pm 0.0009$ ParticleDataGroup:2024cfk . The experimental data are taken from JADE MovillaFernandez:1997fr ; Biebel:1999zt .